# Learning to Evolve: Bayesian-Guided Continual Knowledge Graph Embedding

### v1

@@ Introduction
Figure 3: The change of model performance with the evolution of snapshots under ENTITY and FB_CKGE datasets.
@@
Figure 4: Parameter Sensitivity of BAKE to Cluster Number K under Different Snapshots on the ENTITY Dataset.
Main Results and Analysis. The experimental results are shown in Table 2 and Table 3. Overall, our proposed BAKE achieves the best performance on all datasets, demonstrating the effectiveness of the Bayesian-guided continual learning framework in knowledge graph embedding tasks. First, BAKE achieves the highest MRR scores on all eight datasets, with significant improvements over the strongest baseline methods. Specifically, on the four datasets shown in Table 2, BAKE improves MRR by 10.3% (ENTITY), 4.0% (RELATION), 10.8% (FB-CKGE), and 14.7% (WN-CKGE) compared to the next best method (mainly IncDE). Similar improvements are observed in Hits@1, Hits@3, and Hits@10. Notably, on the large-scale initial KG dataset FB-CKGE, BAKE achieves an MRR of 0.257, a 10.8% improvement over IncDE’s 0.232. These results strongly support that our Bayesian-guided framework can effectively suppress catastrophic forgetting while learning new knowledge.
Secondly, BAKE outperforms all CKGE baselines in most cases. Compared to dynamic architecture methods (PNN and CWR), BAKE improves MRR by 1.7%–19.8%, Hits@1 by 1.4%–13.0%, and Hits@10 by 2.3%–34.5%; compared to replay-based methods (GEM, EMR, and DiCGRL), MRR improves by 2.8%–14.7%, Hits@1 by 2.3%–9.5%, and Hits@10 by 3.8%–23.9%; compared to regularization-based methods (SI, EWC, and LKGE), MRR improves by 0.9%–11.0%, Hits@1 by 0.8%–8.3%, and Hits@10 by 0.7%–17.0%. BAKE also exhibits strong robustness and adaptability on datasets with different knowledge growth patterns. On balanced growth datasets (e.g., GraphEqual) and unbalanced ones (e.g., GraphHigher and GraphLower), BAKE consistently delivers optimal performance. Compared to the second-best baselines, MRR improves by 9.8%, 6.2%, and 8.3%, respectively, on these three datasets. This indicates that our proposed continual clustering constraint mechanism can effectively maintain semantic consistency in the embedding space, regardless of the evolution pattern of the KG.
Finally, comparing our method with baselines from different technical paradigms, BAKE consistently shows superior performance. Whether against dynamic architecture-based methods (PNN, CWR), replay-based methods (GEM, EMR, DiCGRL), or regularization-based methods (SI, EWC, LKGE), BAKE achieves the best results across all metrics. This strongly suggests that framing continual KGE as a sequential Bayesian inference problem, enhanced with explicit semantic consistency constraints, is a more fundamental and effective approach than existing solutions.
Ablation Results and Analysis. In ablation studies conducted on the ENTITY, RELATION, FB-CKGE, and WN-CKGE datasets, removing the Bayesian (Use the distillation ideas from IncDE) resulted in significant overall performance degradation, particularly in Hits@10 and MRR metrics. This highlights its crucial role in preserving prior knowledge through sequential posterior updates and mitigating catastrophic forgetting across evolving snapshots. By contrast, removing the continual clustering mechanism led to an even steeper decline in Hits@1 accuracy. This outcome demonstrates its effectiveness in maintaining semantic consistency and constraining embedding drift for precise link prediction. The integration of these components within BAKE achieves the optimal balance, showcasing synergistic benefits in adaptability and knowledge retention.
Case 1: ($Holland\ Taylor,\ profession,\ ?$)
@@
Case 2: ($Blade\ Runner,\ genre,\ ?$)
@@
Case 3: ($Netherlands,\ contains,\ ?$)
@@
Case 4: ($Raoul\ Walsh,\ place\_of\_birth,\ ?$)
@@
Table 5: Case study experiments on the ENTITY dataset.
Results and analysis on different snapshots. Figure 3 plots the MRR trends of BAKE across 5 snapshots on the ENTITY and FB‑CKGE benchmark datasets. As shown, BAKE consistently outperforms the baselines, achieving the highest and most stable metric values across all snapshots. Examining the curve slopes, BAKE exhibits the flattest decline in performance, indicating its superior resistance to forgetting compared to IncDE and FastKGE. In the continual learning scenario, this capability is demonstrated by BAKE’s Bayesian ”Posterior-as-Prior” update strategy and the semantic cluster consistency regularization via continual clustering. The former facilitates efficient knowledge acquisition, while the latter constrains embedding drift, collectively achieving an optimal trade-off between plasticity (adaptation to new facts) and stability (retention of old knowledge).
The parameter sensitivity experiment of cluster number K. In a Continual learning scenario, traditional K-means clustering directly applied to each snapshot may disrupt stable cluster formation due to the dynamic emergence of new entities. To address this, we fix the cluster size K as a hyperparameter. As shown in Figure 4, the MRR performance of the BAKE model seems to improve with larger K values, indicating that a relatively higher number of clusters enhances semantic consistency and knowledge retention by allowing finer-grained embedding grouping. The performance gap between different K values is most significant in early snapshots and gradually converges as the model evolves over multiple time steps. However, if K is too large, performance may degrade despite increased overhead.
@@
Case Study. To visually demonstrate BAKE’s practical advantages, we conducted an in-depth analysis of four link prediction cases from the ENTITY dataset, comparing BAKE against strong baselines (IncDE and FastKGE). As shown in Table 5, these cases cover diverse relationship types, including profession, category, geographical affiliation, and birthplace. For queries such as (Blade Runner, genre, ?) and (Netherlands, contains, ?), BAKE correctly ranked the ground-truth answers (”Drama” and ”Alkmaar”) at the top position. In contrast, competing models degraded the true answers to positions beyond the top-100. Remarkably, even for highly sparse relations like the query for Raoul Walsh’s birthplace, BAKE achieved a competitive rank (3rd position), significantly outperforming the baselines (179th and 336th ranks).
Conclusion
@@
In this paper, we propose a novel Bayesian-guided CKGE model that effectively mitigates catastrophic forgetting in evolving knowledge graphs by treating sequential updates as Bayesian inference and incorporating continual clustering to preserve semantic consistency. Extensive experiments on multiple benchmarks demonstrate that our model outperforms existing baselines in terms of knowledge preservation and adaptability.

### v2

@@ 5.2. Results and Analysis
Figure 3. The change of model performance with the evolution of snapshots under ENTITY and FB_CKGE datasets.
@@
Figure 4. Parameter Sensitivity of BAKE to Cluster Number K under Different Snapshots on the ENTITY, GRAPH_LOWER, FB_CKGE, and WN_CKGE Datasets.
Main Results and Analysis.
Experimental results are shown in the Table 1 and Table 2. Overall, our proposed BAKE method achieves state-of-the-art performance across all benchmark datasets and evaluation metrics (MRR, Hits@1/3/10), validating the effectiveness of the Bayesian-guided continual learning framework for the continual Knowledge Graph Embedding (CKGE) task. For large-scale initial knowledge graphs (e.g., FB-CKGE) and sparse or difficult scenarios (e.g., WN-CKGE), BAKE achieves state-of-the-art MRR and consistently maintains strong Hits, demonstrating robust knowledge acquisition while resisting catastrophic forgetting. These advantages are consistent with the design goal of combining sequential Bayesian updates with representation-level constraints.
Compared to representative baseline methods from different paradigms, including dynamic architecture methods (PNN, CWR), replay-based methods (GEM, EMR, DiCGRL), regularization-based methods (SI, EWC, LKGE), incremental/distillation or adapter-based methods (IncDE, FastKGE), and recent energy-based/token-driven methods (CLKGE, ETT-CKGE, SAGE), BAKE ranks first on the vast majority of datasets and metrics listed in the Tables 1 and 2, and is at least competitive in the remaining cases. This trend holds across datasets with different growth dynamics: BAKE maintains optimal or near-optimal results for both balanced (GraphEqual) and non-balanced (GraphHigher/GraphLower) evolution modes, demonstrating that the proposed continual clustering constraint effectively stabilizes the embedding geometry under various evolution modes.
From a temporal perspective, BAKE also demonstrates the most stable performance across time snapshots: as the number of snapshots increases, its MRR and hits decay curves are consistently flatter than those of strong baselines such as IncDE/FastKGE, demonstrating its optimal balance between plasticity and stability. In summary, transforming the continual knowledge graph embedding problem into sequential Bayesian inference with explicit semantic consistency constraints provides a more principled and generally effective approach to existing strategies based on heuristic regularization, replay, or architectural expansion.
Ablation Results and Analysis. As shown in Table 3, we conducted ablation experiments on four datasets, demonstrating that our proposed method significantly impacts performance. First, removing the Bayesian (i.e., directly using the IncDE distillation method) significantly degrades the overall model performance: MRR decreases by approximately 7.05% on average, and Hits@10 decreases by approximately 5.84% (e.g., MRR decreases by 8.17% on FB-CKGE and H@10 decreases by 5.37% on ENTITY). This demonstrates that sequential posterior updates effectively preserve prior knowledge and mitigate catastrophic forgetting caused by snapshot evolution. In contrast, removing Continual Clustering leads to steeper fluctuations in H@1: an average decrease of approximately 11.16% (12.74% and 10.32% on FB-CKGE and RELATION, respectively), while H@10 only decreases slightly (1.66% on average). This indicates that this mechanism primarily improves the strict first-place hit rate by maintaining semantic consistency and suppressing embedding drift. Overall, the collaboration of the two modules in BAKE enables it to achieve the optimal trade-off between adaptability and knowledge retention: the Bayesian is responsible for cross-snapshot steady-state memory, and continual clustering finely constrains local discrimination capabilities, thereby benefiting from MRR, Hits@10, and Hits@1 simultaneously.
Figure 5. Knowledge transfer experiments on eight datasets. BAKE w/o B represents the use of knowledge distillation to retain the original knowledge without the guidance of Bayesian principles, and BAKE w/o C represents the use of no continual clustering method.
Table 4. Case study experiments on the ENTITY dataset.
Case 1: (HollandTaylor,profession,?Holland\ Taylor,\ profession,\ ?)
@@
Case 2: (BladeRunner,genre,?Blade\ Runner,\ genre,\ ?)
@@
Case 3: (Netherlands,contains,?Netherlands,\ contains,\ ?)
@@
Case 4: (RaoulWalsh,place_of_birth,?Raoul\ Walsh,\ place\_of\_birth,\ ?)
@@
Results and analysis on different snapshots. Figure  3 shows the MRR trends of BAKE over five consecutive snapshots on the ENTITY and FB-CKGE datasets. Across all time steps, BAKE consistently outperforms strong baseline methods and exhibits the most stable performance trajectory: its MRR decreases more gradually with increasing snapshot index, and its fluctuations across snapshots are smaller. From a mechanistic perspective, the Bayesian approach’s ”posterior is prior” update strategy supports the rapid assimilation of new facts, while the proposed semantic clustering consistency regularization effectively suppresses embedding drift. These two components work together to achieve a favorable trade-off between plasticity and stability, rapidly adapting to new triples while minimizing the loss of existing knowledge.
The parameter sensitivity experiment of cluster number K. In a Continual learning scenario, using traditional K-means clustering independently at each snapshot can lead to cluster drift due to the dynamic emergence of entities and relations. To address this, we set K as a fixed hyperparameter and implement a continual clustering strategy across snapshots. As shown in the Figure 4, when K increases within a reasonable range, the MRR of BAKE improves overall, demonstrating that finer-grained clustering enhances semantic consistency and knowledge retention by providing a more accurate neighborhood structure. The performance differences between different K values are most pronounced in early snapshots and gradually decrease as the model evolves, indicating that the sensitivity to the number of clusters decreases after a stable representation is formed. However, excessively large K values can lead to semantic fragmentation and introduce additional overhead, resulting in diminishing returns or even negative effects. In practical applications, choosing a moderate K value can ensure robust performance while avoiding instability and unnecessary computational cost.
@@
Case Study. To visually demonstrate BAKE’s practical advantages, we conducted an in-depth analysis of four link prediction cases from the ENTITY dataset, comparing BAKE against strong baselines (IncDE and FastKGE). As shown in Table 4, these cases cover diverse relationship types, including profession, category, geographical affiliation, and birthplace. For queries such as (Blade Runner, genre, ?) and (Netherlands, contains, ?), BAKE correctly ranked the ground-truth answers (”Drama” and ”Alkmaar”) at the top position. In contrast, competing models degraded the true answers to positions beyond the top 100. Remarkably, even for highly sparse relations like the query for Raoul Walsh’s birthplace, BAKE achieved a competitive rank (3rd position), significantly outperforming the baselines (179th and 336th ranks).
Knowledge Transfer Capability. We evaluate knowledge transfer using the FWT metric on eight datasets. The experimental results are shown in Figure 5. BAKE attains the best FWT on all datasets, significantly outperforming IncDE (incremental distillation), GEM (replay-based), and both abridged variants. The gain is larger on datasets with more dramatic structural shifts (e.g., GraphHigher, GraphLower, FB-CKGE), showing that BAKE transfers learned concepts more effectively to later snapshots. Removing the Bayesian posterior update (BAKE w/o B) or the continual clustering constraint (BAKE w/o C) reduces FWT, indicating these components jointly suppress representation drift and improve forward transfer.
6. Conclusion
@@
In this paper, we propose a novel Bayesian-guided CKGE model that effectively mitigates catastrophic forgetting in evolving knowledge graphs. Our framework comprises two complementary modules: (1) Bayesian-guided knowledge evolution, which treats sequential updates as Bayesian inference and uses posterior-as-prior updates to provide theoretical guarantees for knowledge preservation, and (2) continual clustering, which constrains semantic drift across snapshots. This ensures entities remember not only ”who they are” (intra-cluster compactness) but also ”who they are related to” (inter-cluster separability). Extensive experiments on multiple benchmarks demonstrate that our model significantly outperforms existing baselines in both knowledge preservation and adaptability, achieving optimal balance between plasticity and stability.
Appendix A Appendix
A.1. Detailed Experiment Settings
A.1.1. Datasets.
ENTITY (Cui et al., 2023): This dataset simulates a growth scenario dominated by newly added entities. New entities are uniformly introduced and triple connections are supplemented at each time slice, requiring the model to learn new entity representations while preserving existing knowledge. This dataset contains five snapshots and serves as a benchmark for ”new entity-driven evolution” in continual learning.
RELATION (Cui et al., 2023): This model depicts growth driven primarily by newly added relationships. New relationship types are added evenly over time, with the number of triples growing fastest. This model is used to test the stability of the model under frequent relationship additions. Five snapshots are also provided.
HYBRID (Cui et al., 2023): This model simulates ”non-uniform dimensional growth,” which is closer to reality. Entities, relationships, and triples all evolve, resulting in large incremental fluctuations and a higher level of challenge, which better tests the robustness of the method. Five snapshots are also provided.
GraphEqual/Higher/Lower (Liu et al., 2024a) all remove the restriction that new triples must contain old entities, to more closely resemble real-world graph evolution. They differ in their growth patterns: GraphEqual’s increments are equal across all time periods, assessing continual learning under balanced growth; GraphHigher’s increments increase gradually over time, simulating an accelerated burst of knowledge and testing the model’s performance under sustained pressure; and GraphLower’s increments are large early on and decreasing later, testing the model’s ability to resist forgetting and consolidate knowledge during a ”fast-first-then-slow” evolution.
FB-CKGE (Liu et al., 2024b) and WN-CKGE (Liu et al., 2024b) are two benchmarks built on FB15k-237 and WN18RR. They allocate about 60% of triples in the initial snapshot, forming a large initial graph with a cold start, then gradually add the rest. This design tests methods for efficiency and anti-forgetfulness in large initial graphs with incremental updates.
A.1.2. Baselines.
Fin-tune (direct fine-tuning) (Cui et al., 2023) trains only on new triplets and often forgets old knowledge; Elastic Weight Consolidation (EWC) adds penalties on parameters important to past tasks to preserve them; PNN (Progressive Neural Networks) (Rusu et al., 2016) adds a new column per task with lateral connections to avoid forgetting, trading off with parameter growth.
IncDE(Incremental Distillation) (Liu et al., 2024a) . Uses hierarchical learning on new triplets with layer-wise distillation; a two-stage routine reduces early noise and better retains old knowledge.
FastKGE (Incremental LoRA) (Liu et al., 2024b). Inserts LoRA adapters only where affected, adaptively sets ranks to cut parameters/training time, and uses layer selection to limit forgetting.
LKGE (Lifelong KGE) (Cui et al., 2023) Updates embeddings via a masked KG autoencoder and uses transfer + regularization so new entities/relations inherit useful prior knowledge.
SI (Synaptic Intelligence) (Zenke et al., 2017). Online tracks each parameter’s contribution to loss reduction, assigns importance, and regularizes updates to protect critical weights.
DiCGRL (Disentangle-based Continual Graph Representation Learning (Kou et al., 2020)). Separates transferable from task-specific factors in graph representations to learn new data while reducing forgetting.
EMR (Episodic Memory Replay) (Wang et al., 2019) replays a small buffer of old samples with new ones (EA-EMR aligns embeddings to reduce drift); CWR (Copy Weights with Re-init) (Lopez-Paz and Ranzato, 2017) resets a temporary head per task and copies it to a consolidated head to limit interference; FMR (Flexible Memory Rotation) (Zhu et al., 2024) rotates updates in parameter space (guided by Fisher information) and tunes regularization strength to balance stability and plasticity.
ETT-CKGE (Efficient Task-driven Tokens for CKGE (Zhu et al., 2025)). Replaces explicit scoring/traversal with task-driven learnable tokens for alignment/transfer across snapshots, enabling simple matrix ops and lower training cost.
SAGE (Scale-Aware Gradual Evolution(Li et al., 2025b)). Adapts embedding size to update scale and uses dynamic distillation to balance retention and absorption, yielding robust gains across benchmarks.
CLKGE  (Cao, 2024). Learns an energy manifold for dynamic KGs and aligns old/new energies to mitigate forgetting while transferring knowledge forward.
A.1.3. FWT Metric.
To measure forward transferability, that is, the immediate help of the knowledge learned in the previous snapshot for the new task in the next snapshot, we use FWT to measure the ability of the model:
(10) $\mathrm{FWT}=\frac{1}{N}\sum_{i=1}^{N}a_{i-1,i},$
where $a_{t,i}$ represents the score of the model $M_{t}$ trained on the $t$th snapshot on the test set (often expressed as MRR or Hits@K) at the $i$th snapshot. Therefore, $a_{i-1,i}$ captures the model’s zero-shot generalization ability to the next task before seeing $\Delta T_{i}$ training data. A larger FWT value indicates that the method is more effective in transferring learned structural and semantic patterns to subsequent snapshots.
Table 5. The statistics of datasets. $N_{E}$, $N_{R}$ and $N_{T}$ denote the number of entities and relations, and the current snapshot $i$
Snapshot 0 Snapshot 1 Snapshot 2 Snapshot 3 Snapshot 4
Dataset $N_{E}$ $N_{R}$ $N_{T}$ $N_{E}$ $N_{R}$ $N_{T}$ $N_{E}$ $N_{R}$ $N_{T}$ $N_{E}$ $N_{R}$ $N_{T}$ $N_{E}$ $N_{R}$ $N_{T}$
A.2. Related Work on Continual Learning
Continual learning (CL) enables models to learn from a continuous data stream, addressing the ”catastrophic forgetting” of old knowledge, a problem first identified in neural networks by McCloskey and Cohen(McCloskey and Cohen, 1989). The goal is to balance stability (retaining knowledge) with plasticity.
CL strategies are now applied across various dynamic environments. Recent examples include Helios (Shi et al., 2025) for adapting to new cyber threats, DSLR (Choi et al., 2024) for enhancing diversity in graph continual learning, and Multimodal Graph Continual Learning (Cai et al., 2022) over evolving document corpora. In recommender systems, Continual Collaborative Distillation (Lee et al., 2024a) allows models to adapt to non-stationary user and item data. These advancements demonstrate CL’s growing importance in web systems.
A.3. BAKE model training process pseudocode
To help readers more clearly and comprehensively understand our proposed BAKE model, we present its training process in detail using pseudocode. The core idea of BAKE is to formulate the CKGE task as a sequential Bayesian inference problem, thus providing a more theoretically complete framework for the continuous accumulation of knowledge and the suppression of forgetting. The algorithm iterates through each knowledge graph snapshot sequentially. For a given snapshot $t$, the model utilizes the parameterized posterior distribution $(\boldsymbol{\mu}_{t-1},\boldsymbol{\lambda}_{t-1})$ of the previous snapshot t1t-1 as a guiding prior. During training, the model jointly optimizes three objectives: a standard KGE loss (for learning new knowledge), a Bayesian regularization term guided by the prior (for preventing catastrophic forgetting of old knowledge), and a persistent clustering loss (for maintaining semantic consistency in the embedding space and preventing representation drift). After training on the current snapshot data, the model performs a Bayesian update to compute a new posterior distribution $(\boldsymbol{\mu}_{t},\boldsymbol{\lambda}_{t})$, which serves as the prior for the next snapshot $t+1$. Algorithm 1 describes in detail the complete steps from initialization to the final sequential update.
Algorithm 1 BAKE Model Training Process
1:KG snapshot sequence $\mathcal{G}_{c}=\{\mathcal{S}_{0},\dots,\mathcal{S}_{N}\}$; hyperparameters: lr, epochs $E$, reg. $\beta$, obs. precision $\lambda_{obs}$, momentum $\eta$, clusters $K$, temp. $\tau$.
2:Init for $\mathcal{S}_{0}$: init entity/relation embeddings $\hat{\boldsymbol{\Theta}}_{0}$; set priors $\boldsymbol{\mu}_{0}\!\leftarrow\!\hat{\boldsymbol{\Theta}}_{0}$, small $\boldsymbol{\lambda}_{0}$; compute $IE(e)$ and cluster into $K$; get centroids $\mathbf{c}_{k,0}$ and proxies $\mathbf{v}_{k,0}$.
3:for $t=0\to N$ do
4:  if $t>0$ then
5:   Inherit priors $\{\boldsymbol{\mu}_{t-1},\boldsymbol{\lambda}_{t-1}\}$; init new entities $\Delta\mathcal{E}_{t}$; set $\mathbf{c}_{k,t}\!\leftarrow\!\mathbf{c}_{k,t-1}$, $\mathbf{v}_{k,t}\!\leftarrow\!\mathbf{v}_{k,t-1}$.
6:  end if
7:  for $epoch=1\to E$ do $\triangleright$ Train on $\mathcal{S}_{t}$
8:   for each batch in $\mathcal{T}_{t}$ do
9:     Compute KGE loss $\mathcal{L}_{\text{KGE}}$ via eq. 1.
10:     Compute Bayesian reg. $\mathcal{L}_{\text{Bayes}}$ via eq. 4 using $\{\boldsymbol{\mu}_{t-1},\boldsymbol{\lambda}_{t-1}\}$.
11:     Reassign entities to nearest clusters; compute $\mathcal{L}_{\text{FCC}}$ via eq. 6.
12:     $\mathcal{L}_{\text{total}}\leftarrow\mathcal{L}_{\text{KGE}}+\mathcal{L}_{\text{Bayes}}+\mathcal{L}_{\text{FCC}}$; update $\hat{\boldsymbol{\Theta}}_{t}$, $\mathbf{v}_{k,t}$.
13:   end for
14:   Update centroids $\mathbf{c}_{k,t}$ with momentum via eq. 8.
15:  end for
16:  Bayesian update for next snapshot:
17:  $\boldsymbol{\lambda}_{t}\leftarrow\boldsymbol{\lambda}_{t-1}+\lambda_{obs}$ (via eq. 2)
18:  $\boldsymbol{\mu}_{t}\leftarrow\frac{\boldsymbol{\lambda}_{t-1}\odot\boldsymbol{\mu}_{t-1}+\lambda_{obs}\odot\hat{\boldsymbol{\Theta}}_{t}}{\boldsymbol{\lambda}_{t}}$ (via eq. 3)
19:end for
20:return final embeddings $\hat{\boldsymbol{\Theta}}_{N}$.